# Filtered Corpus Training (FiCT) Shows that Language Models can Generalize from Indirect Evidence

### v1

@@ 7 Conclusion
We introduced the Filtered Corpus Training methodology and applied it to a wide range of linguistic constructions from the BLiMP benchmark. Our results show that while Transformers are better language models (via perplexity) than comparable LSTMs, the latter generalize equally well (via $\textsf{acc}\Delta$ and $P\Delta$). The relatively low $\textsf{acc}\Delta$ scores in general show that all of our LMs exhibit a strong ability to generalize from indirect evidence, even for models of relatively low parameter count trained on relatively small data. Future work will (i) extend this approach to models of different sizes and pretraining corpora, (ii) perform deeper analyses of the bases on which the models do make their generalizations, and (iii) analyze other forms of lexical and structural generalization through the lens of filtered corpus training.
@@ Appendix A Training Hyperparameters
Table 2: Selected training hyperparemeters, as provided to the transformers package’s TrainingArguments class. Any omitted values were set to the defaults associated with version 4.30.2 of the transformers package.

### v2

@@ 7 Conclusion
We introduced the Filtered Corpus Training methodology and applied it to a wide range of linguistic constructions from the BLiMP benchmark. Our results show that while Transformers are better language models (via perplexity) than comparable LSTMs, the latter generalize equally well (via $\textsf{acc}\Delta$ and $P\Delta$). The relatively low $\textsf{acc}\Delta$ scores in general show that all of our LMs exhibit a strong ability to generalize from indirect evidence, even for models of relatively low parameter count trained on relatively small data. In summary, this shows that language model success cannot be attributed solely to memorization from its training data, since the data has been systematically purged of the evaluation targets. They are, instead, able to form subtle and linguistically-relevant generalizations from indirect evidence.
Future work will (i) extend this approach to models of different sizes and pretraining corpora, (ii) perform deeper analyses of the bases on which the models do make their generalizations (including with probing experiments), and (iii) analyze other forms of lexical and structural generalization through the lens of the filtered corpus training methodology.
@@ Appendix A Training Hyperparameters
Table 2: Selected training hyperparameters, as provided to the transformers package’s TrainingArguments class. Any omitted values were set to the defaults associated with version 4.30.2 of the transformers package.